# An engineering guide to superconducting quantum circuit shielding

### v1

@@ I INTRODUCTION
One of the most promising physical realizations of quantum computers and simulators are superconducting quantum circuits (SQC) Nakamura, Pashkin, and Tsai (1999); Blais et al. (2021); Krantz et al. (2019); Kjaergaard et al. (2020). Nowadays, there are several superconducting quantum computer prototypes capable to solve very complex specific tasks, which are not amenable for the state-of-the-art supercomputers Arute et al. (2019); Morvan et al. (2024); Acharya et al. (2025); Gao et al. (2025); Kim et al. (2023). However, quantum information loss or decoherence is still one of the key challenges on the way to the useful quantum computers. Robust error-free operation of the superconducting qubits, which are extremely sensitive to noise Bylander et al. (2011); Koch, DiVincenzo, and Clarke (2007); Yan et al. (2012); Bialczak et al. (2007); Schuster et al. (2005), requires minimizing both external and internal sources of noise. Internal sources can be combated by improving the design Ganjam et al. (2024); Pan et al. (2022), materials Kim et al. (2021); Place et al. (2021); Oliver and Welander (2013) and technology Pishchimova et al. (2023); Moskalev et al. (2023) of quantum circuits.
In the last two decades, the leading scientific groups in the field demonstrated tremendous results in minimizing the internal noise sources including magnetic vortices Ithier et al. (2005), quasiparticle tunneling Córcoles et al. (2011); Zaretskey et al. (2013); de Visser et al. (2011), two-level systems Ithier et al. (2005); Zaretskey et al. (2013) and dielectric losses in substrate Calusine et al. (2018); Wang et al. (2015); Woods et al. (2019). A number of papers are devoted to an external noise sources suppression, such as flux and charge noise Ithier et al. (2005), Purcell effect Purcell (1946), interaction with stray electromagnetic modes of a cryogenic experimental setup or environment Córcoles et al. (2011); Zaretskey et al. (2013). However, superconducting circuits protection from a broadband spontaneous radiation: from surrounding equipment (frequency below 300 $GHz$, wave length more 1 $mm$), infrared (IR) radiation (frequency from 300 $GHz$ to 430 $THz$, wave length from 0.74 ${\mu}m$ to 1 $mm$), cosmic rays, and background radiation (frequency above 430 $THz$, wave length less 0.74 ${\mu}m$) Córcoles et al. (2011); Martinis (2021) is currently being researched.
@@
Articles by leading scientific groups show a wide variety of shielding systems, which differ greatly in the number of shields and their combination Eddins et al. (2018); Burnett et al. (2019); Schwartz (2016); Place et al. (2021); Premkumar et al. (2021); Yan et al. (2016); Schmitt (2015); Burnett (2013); Córcoles et al. (2015). Comparing represented shielding systems, it can be seen that they differ from each other by the number of shields, their relative position, position inside the dilution refrigerator, and the design and materials used.
@@
    Versatility of shielding. Considered shielding systems consist of shielding layers protecting from both IR radiation and the magnetic field. Additionally, in some cases a single shield can combine the functions of IR and magnetic shielding. For this purpose, an IR-absorbent coating is applied to either a superconducting or magnetic metal shield.
@@
    IR shielding implementation. IR shielding is composed of a metal base with a specialized IR-absorbing coating based on the epoxy resin or mixture of epoxy resin and silicon carbide particles or graphite powder. This coating should exhibit a low level of reflectivity (up to 0.0001), i.e., high IR absorption.
@@
The filtering systems are also varied from one scientific group to another. The filters are an integral part of the cryogenic setup for quantum systems. They are employed in all control and measurement lines of the fridge, specifically in readout, drive, and flux lines. There are many options of the filters placement in cryogenic measurement setup (see Fig. 1). In the same way as for the shielding systems, choosing the optimal filter position is also a challenging task.
This review represents set of requirements for choosing shielding and filtering system for effective protection of SQC from IR radiation and unwanted electromagnetic fields based on the analysis of existing multilayer shielding designs, used materials and signal lines filtering and attenuation.
@@ I.1 Article organization
In the Section II, we discuss what type of shielding is required for protection from infrared radiation and stray electromagnetic fields. We present the requirements for an effective shielding system based on theoretical estimations and numerical modeling, and analyze existing shielding systems according to these requirements. We propose additional shielding techniques to improve protection efficiency, and demonstrate how the sample holder design affects the shielding effectiveness.
@@
In summary and conclusions, we discuss the similarities and differences in the shielding of quantum circuits, and we also attempt presenting our recommendations for the compact and efficient quantum circuit protection against external electromagnetic fields.
@@ II.1.1 Shielding principle
Protection from an IR radiation is based on the absorption of the energy from incident photons by the absorbing material of a shield and the conversion of this energy into internal thermal energy of the shield. This thermal energy is then removed by the cooling mechanisms within the fridge. As a result, the photons partially or completely lose the energy and can’t cause Cooper pairs braking in superconductor. To protect from the described source of an IR radiation, the shields and, in some cases, sample holder lids are coated by special IR absorbing materials. The absorption coefficient of such coatings can reach the value up to 0.9 in THz range of the electromagnetic spectrum Barends et al. (2011). Commercially available epoxy resins, such as Stycast 2850 FT and Eccosorb CR series, are used as IR-absorbing coatings, either alone or in combination with other particles such as silicon carbide powder, carbon powder, or graphite dust.
@@ II.1.2 Requirements for IR shielding
Based on the theory of heat transfer by radiation, the following concepts for an IR shielding can be formulated:
@@
    To improve IR radiation protection (see simulation below) and reduce weight we recommend the inner surface of the shield with an absorption coefficient close to 1 and the outer surface – approaching 0 (highly reflective) in order to reflect outside radiation.
@@
    Shield outer surface can be made of a smooth and polished metal to maximize reflection of outside IR radiation (minimal absorption coefficient).
@@
    In order to increase the absorption coatings capacity one should insure its surface roughness on the order of the IR wavelength Smith (1984). This level of roughness avoid specular reflection, scattering the incident IR radiation, which undergoes multiple reflections and corresponding interactions within absorption material, extending an effective surface and overall absorption. In cryogenic quantum applications a far-IR range with wavelengths of 700 – 1000 ${\mu}$$m$ become one of the most important Ekin (2006), consequently, macroscopic roughness of the same size should be introduced for better absorption capacity.
@@
    The temperature of the shield must be lower than the temperature of the quantum circuit to ensure the heat transfer from the sample to the shield. The shield base should be made of a metal with the highest possible thermal conductivity and have robust thermal contact between the shield and the lower stage of the fridge.
Following these requirements is the reliable way for creating an effective IR shielding. Let’s consider the existing IR shields designs.
@@ II.1.3 Practical implementation of IR shielding
Based on the requirements for IR shielding, it is necessary to use a shield base made of material with a high thermal conductivity. The thermal conductivity of some materials commonly used for the base of IR shields is represented in Table 1 below.
@@
[…]
${\mu}$-metal 19 shield Corp. -
It can be clearly seen from Table 1, that high purity copper has a higher thermal conductivity, making it preferable material for the base of IR shielding.
The use of copper for the shielding base is also supported by shielding systems described the literature. Fig. 2 shows the most commonly used materials for the base of IR shielding described in Schwartz (2016); Chou (2018); Dickel et al. (2018); Pop et al. (2014); Monroe et al. (2021a). More than one-third  [^1] Here and further, the estimate of mention frequency of a certain parameter is given from the total number of works where this parameter is found of all listed in Fig. 2 IR shields combine the IR and magnetic shielding. The base in these shielding systems is made of another materials, such as copper coated with a superconductor Premkumar et al. (2021); Henriques et al. (2019), in particular superconducting aluminum Gyenis et al. (2021); Hazard (2019); Smith (2019), or ${\mu}$-metal Poletto et al. (2012); Beck (2018). The combination of IR and magnetic shielding allows for a more compact and lightweight shielding design. However, heat removal capabilities of these shielding systems are reduced compared to the system with a copper base.
Rarely, gold-plated copper Wu et al. (2021); Ann (2021); Chiorescu et al. (2003) is used as the base material for the IR shield. Gold-plating of the entire shielding components does not provide any advantages, as it reduces thermal conductivity at cryogenic temperatures compared to uncoated copper. However, gold-plating is highly beneficial for thermal interfaces between shielding components. Gold-gold contacts demonstrate over an order of magnitude greater thermal conductance than copper-copper contacts Ekin (2006) due to the absence of surface oxidation and improved conformability Dhuley (2019). This enhancement relaxes the required surface finishing quality and contact pressure for copper interfaces.
@@
Another important component of IR shielding is an IR absorbing coating. In reference Bründermann and Hubers (2012), some absorbing coatings data, that are used in the IR radiation range, are provided (Eccosorb CR110, Herberts 1356H, TK RAM, and silicon carbide (SiC) grains in Stycast). Assuming that the IR shielding material does not transmit radiation, we can consider the sum of absorption and reflection coefficients to be 1. This allows us to estimate the absorption properties based on the reflection coefficient. According to theory Howell, Menguc, and Siegel (2010), absorbent coatings should exhibit maximum absorption (near 1) in the operating range. Among the absorbent coatings shown in reference Bründermann and Hubers (2012), SiC grains in Stycast have the highest absorption capacity. For IR shielding, Stycast epoxy-based coatings are commonly used both in a pure form (references Grünhaupt et al. (2017); Maleeva et al. (2018)) and in combination with additional particles to enhance surface roughness (reference Barends et al. (2011)) (see Fig. 3). To form a relief of this surface, SiC particles of various sizes Dickel et al. (2018); Hazard (2019); Jin et al. (2015) or carbon powder Chou (2018); Pop et al. (2014); Wang et al. (2019) or thereof combination Burnett et al. (2019); Ann (2021), can be used. The use of 1 $mm$ size SiC particles mixed with Stycast epoxy is preferrable due to its highest absorption capacity. Such particle size ensures the fulfillment of the surface roughness requirement for the IR range. Such particle size ensures the fulfillment of the surface roughness requirement for the IR range.
@@
Other absorbent materials that may be used include epoxy resins and Eccosorb foams Beck (2018). Eccosorb CR-124 Córcoles et al. (2011); Zhang (2018); Srinivasan (2013); Chow et al. (2014) and Eccosorb MCS in a paste form Liu (2016a) are applied to the metal base. Eccosorb foam Kroll (2019); Rigetti et al. (2012); Uilhoorn (2021) or MCS tape Liu et al. (2014) may be wrapped directly around the sample holder. The thermal conductivity of these absorbent tapes and foams are on the order or less than 1 $W\cdot m^{-1}\cdot K^{-1}$ Laird Technologies , that is much lower in comparison with metals thermal conductivity. At cryogenic temperatures the difference in thermal conductivity between Eccosorb and metals (base materials) increase due to the suppression of phonon thermal conductivity. For example, at around 1 $K$, the difference is already more than 5 orders of magnitude (1000 $W\cdot m^{-1}\cdot K^{-1}$ for oxygen-free copper compared to 0.0038 $W\cdot m^{-1}\cdot K^{-1}$ for Eccosorb Ekin (2006)). With that in mind, it’s recommended to use such absorbent materials in pair a high thermal conductivity metal base for better cooling efficiency.
@@
In addition to IR protection, thin metallic layers are used to reflect spontaneous parasitic IR radiation. For this purpose, metallic foils Schwartz (2016), films Kroll (2019) or aluminized polyethylene terephthalate (trade name Mylar) Place et al. (2021); Gyenis et al. (2021); Rigetti et al. (2012); Narla (2017) may be used. The placement of this layer may vary: it may be placed around the holder Place et al. (2021), around the first shield Hazard (2019); Rigetti et al. (2012), between the shields Premkumar et al. (2021); Gyenis et al. (2021); Hazard (2019) or surround the entire shielding system Place et al. (2021); Premkumar et al. (2021); Hazard (2019). At the same time, the outer surface of the shield may act as a reflective layer in case it has a low absorption coefficient. This allows for the creation of a more compact shielding system.
@@
A more effective shielding system can be defined using the theory of a heat transfer by radiation and modeling the radiation propagation in the "sample-shielding-dilution refrigerator" system. A calculation scheme is represented in Fig. 4. In this system, "Body 1" is the sample, "Body 2" is the dilution refrigerator or radiation source, and "Shields" is the shielding system, which effectiveness we’d like to estimate. The power of a heat transfer by radiation is calculated by the following relation:
@@
In case the shield has different absorption coefficients of its inner and outer surfaces, formula for $A_{(1,2)sh}$ will take: $A_{(1,2)sh}=\frac{1}{\frac{1}{A_{1,2}}+\sum\limits_{i=1}^{n}\frac{F_{1}}{F_{shi}}\left({\frac{1}{A_{shi(inner)}}+\frac{1}{A_{shi(outer)}}-1}\right)};$
@@
Figure 4: Calculation scheme used for the heat transfer (by radiation) calculations between two bodies with an arbitrary number of shielding elements.
@@
Figure 5: Shielding system simulations: symbols and simulated configurations. Shielding configurations: A – copper sample holder with gold-plated copper can around the lowest temperature stage of a fridge (no absorption coatings); B – copper sample holder with the can with absorption coating; C – copper sample holder with an inner absorption coating with the can (no absorption coatings); D – copper sample holder with an outer absorption coating with the can (no absorption coatings); E – copper sample holder with both side absorption coatings and the can; F – sample holder (inner absorption coating) with additional shield around the sample holder (inner absorption coating) and gold-plated copper can; G – sample holder (inner absorption coating) with two additional shields around the sample holder (inner absorption coating) and gold-plated copper can. Symbols: SQC – superconducting quantum circuit; SH – sample holder; FS – first shield; SS – second shield; C – lower stage can of a fridge; HS – heat source.
The results of the calculations is verified using finite element simulations. Computer calculations is performed using COMSOL Multiphysics Heat Transfer in Solids and Surface to Surface Radiation. A two-dimensional axisymmetric model (see Fig. 6) was constructed to simplify calculations. We use the same surface absorption coefficients in the model for all the calculated scheme ($A_{Si}$ = 0.77, $A_{Al}$ = 0.057, $A_{Cu}$ = 0.0134, $A_{Au}$ = 0.0479, $A_{Absorp.coat.}$ = 0.9524). The declared heat dissipation power is $3.45\cdot 10^{-9}W$. The temperature of the lower plate of the cryostat is fixed at 0.01 $K$. Two cases were considered, when the heating source is a superconducting quantum circuit itself (SQC, Fig. 5), and an external heat source (HS, Fig. 5). All the temperatures are normalized relative to the highest temperature, which we have got in all the simulations. The shown simulation result is the final temperature of a superconducting quantum circuit after stabilization for different shielding systems. The shielding system efficiency can also be estimated based on the minimum attainable temperature of the sample.
@@
In Fig. 5 we introduce the following designations: “SQC” – superconducting quantum circuit, “SH” – sample holder, “FS” – first shield, “SS” – second shield, “HS” – heat source in the form of an IR filter, and “C” – can on the lower stage of the fridge.

### v2

@@ I INTRODUCTION
One of the most promising physical realizations of quantum computers and simulators are superconducting quantum circuits (SQC) Nakamura, Pashkin, and Tsai (1999); Blais et al. (2021); Krantz et al. (2019); Kjaergaard et al. (2020). Nowadays, there are several superconducting quantum computer prototypes capable to solve very complex specific tasks, which are not amenable for the state-of-the-art supercomputers Arute et al. (2019); Morvan et al. (2024); Acharya et al. (2025); Gao et al. (2025); Kim et al. (2023). However, quantum information loss or decoherence is still one of the key challenges on the way to the practically useful quantum computers. Robust error-free operation of the superconducting qubits, which are extremely sensitive to noise Bylander et al. (2011); Koch, DiVincenzo, and Clarke (2007); Yan et al. (2012); Bialczak et al. (2007); Schuster et al. (2005), requires minimizing both external and internal sources of noise. Internal sources can be suppressed by improving the design Ganjam et al. (2024); Pan et al. (2022), materials Kim et al. (2021); Place et al. (2021); Oliver and Welander (2013) and technology Pishchimova et al. (2023); Moskalev et al. (2023) of quantum circuits.
In the last two decades, the leading scientific groups in the field demonstrated tremendous results in minimizing the internal noise sources including magnetic vortices Ithier et al. (2005), quasiparticle tunneling Córcoles et al. (2011); Zaretskey et al. (2013); de Visser et al. (2011), two-level systems Ithier et al. (2005); Zaretskey et al. (2013) and dielectric losses in substrate Calusine et al. (2018); Wang et al. (2015); Woods et al. (2019). A number of papers are devoted to an external noise sources suppression, such as flux and charge noise Ithier et al. (2005), Purcell effect Purcell (1946), interaction with stray electromagnetic modes of a cryogenic experimental setup or environment Córcoles et al. (2011); Zaretskey et al. (2013). However, superconducting circuits protection from a broadband spontaneous radiation is currently being researched: from surrounding equipment (frequency below 300 $GHz$, wave length more 1 $mm$), infrared (IR) radiation (frequency from 300 $GHz$ to 430 $THz$, wave length from 0.74 ${\mu}m$ to 1 $mm$), cosmic rays, and background radiation (frequency above 430 $THz$, wave length less 0.74 ${\mu}m$) Córcoles et al. (2011); Martinis (2021) is currently being researched.
@@
Articles by leading scientific groups show a wide variety of shielding systems, which differ greatly in the number of shields and their combination Place et al. (2021); Burnett et al. (2019); Kreikebaum et al. (2016); Yan et al. (2016); Grünhaupt et al. (2018); Bengtsson (2020); Reagor (2015); Asaad et al. (2016); Janzen et al. (2023). Represented shielding systems comparison shows that they differ from each other by the number of shields, their relative position, position inside the dilution refrigerator, and the design and materials used.
@@
    Shielding versatility. The considered shielding systems comprise multiple layers providing protection against both IR radiation and magnetic fields. In certain configurations, a single shield can simultaneously serve as both IR and magnetic shielding. This dual functionality is achieved by applying an IR-absorbent coating to a superconducting shield.
@@
    IR shielding implementation. The IR shielding consists of a metallic substrate coated with a specialized IR-absorbing material. This coating is typically produced using epoxy resins, either pure or mixed with silicon carbide (SiC) particles or graphite powder. The primary requirement for these coatings is achieving high IR absorption efficiency.
@@
The filtering systems are also varied from one scientific group to another. The filters are an integral part of the cryogenic setup for quantum systems. They are employed in all control and measurement lines of the fridge, specifically in readout, drive, and flux lines. There are many options of the filters placement in cryogenic measurement setup (see Figure 1). In the same way as for the shielding systems, choosing the optimal filter position is also a challenging task.
This review represents set of requirements for choosing shielding and filtering system for effective protection of SQC from IR radiation and unwanted electromagnetic fields based on the analysis of existing multilayer shielding designs, used materials and signal lines filtering.
@@ I.1 Article organization
In the Section II, we discuss what type of shielding is required for protection from infrared radiation and stray electromagnetic fields. We present the requirements for an effective shielding system based on theoretical estimations and numerical modeling, and analyze existing shielding systems according to these requirements.
@@
In summary and conclusions we consolidate our recommendations for fabricating compact and efficient shielding systems to protect quantum circuits from external electromagnetic fields.
@@ II.1.1 Shielding principle
Protection from an IR radiation is based on the absorption of the energy from incident photons by the absorbing material of a shield and the conversion of this energy into internal thermal energy of the shield. This thermal energy is then removed by the cooling mechanisms within the fridge. As a result, the photons partially or completely lose the energy and can’t cause Cooper pairs braking in superconductor. To protect from the described source of an IR radiation, the shields and, in some cases, sample holder lids are coated by special IR absorbing materials. Commercially available epoxy resins, such as Stycast 2850 FT and Eccosorb CR series, are used as IR-absorbing coatings, either alone or in combination with other particles such as silicon carbide powder, carbon powder, or graphite dust.
@@ II.1.2 Requirements for IR shielding
Guided by radiative heat transfer theory, we derive the core requirements for effective IR shielding of quantum devices:
@@
    Optimal IR radiation shielding requires configuring the shield’s inner surface with an absorption coefficient ($A$) approaching 1 to effectively absorb penetrating radiation, while the outer surface should maintain $A{\approx}0$ to maximize reflection of incident external radiation.
@@
    To maximize reflection of incident IR radiation, the shield’s outer surface should be fabricated from a polished, highly reflective metal.
@@
    To increase the absorption coatings capacity one should insure its surface roughness on the order of the IR wavelength Smith (1984). This level of roughness avoid specular reflection, scattering the incident IR radiation, which undergoes multiple reflections and corresponding interactions within absorption material, extending an effective surface and overall absorption. In cryogenic quantum applications a far-IR range with wavelengths of 700 – 1000 ${\mu}$$m$ become one of the most important Ekin (2006), consequently, macroscopic roughness of the same size should be introduced for better absorption capacity.
@@
    The temperature of the shield must be lower than the temperature of the quantum circuit to ensure the heat transfer from the sample to the shield. The shield base should be made of a metal with the highest possible thermal conductivity at operating temperature and have robust thermal contact between the shield and the lower stage of the fridge.
Following these requirements is the proper way for creating an effective IR shielding. Let’s consider the existing IR shields designs.
@@ II.1.3 Practical implementation of IR shielding
The requirements for IR shielding shows, that it is necessary to use a shield base made of material with a high thermal conductivity. The thermal conductivity of some materials commonly used for the base of IR shields is represented in Table 1 below.
As shown in Table 1, high purity copper has a higher thermal conductivity, making it preferable material for the base of IR shielding.
The use of copper for the shielding base is also supported by shielding systems described in the literature. Figure 2 shows the most commonly used materials for the base of IR shielding described in Schwartz (2016); Chou (2018); Dickel et al. (2018); Monroe et al. (2021a). Approximately 65%  [^1] Here and further, the estimate of mention frequency of a certain parameter is given from the total number of works where this parameter is found of shielding configurations (Figure 2) integrate combined IR/magnetic shielding, using either copper substrates with superconducting coatings (aluminum (Al) Grünhaupt et al. (2018); Henriques et al. (2019), indium (In) Premkumar et al. (2021), or tin (Sn) Kreikebaum et al. (2016)) or bulk superconductors Smith (2019); Zhang (2018); Gyenis et al. (2021); Hazard (2019). This combined shielding approach enables compact and lightweight shielding designs. However, pure superconducting shields exhibit significantly lower heat dissipation capacity than their copper-based counterparts.
@@
Rarely, gold-plated copper Byoung-moo (2021); Wu et al. (2021) is used as the base material for the IR shield. Gold-plating of the entire shielding components does not provide any advantages, as it reduces thermal conductivity at cryogenic temperatures compared to uncoated copper. However, gold-plating is highly beneficial for thermal interfaces between shielding components. Gold-gold contacts demonstrate over an order of magnitude greater thermal conductance than copper-copper contacts Ekin (2006) due to the absence of surface oxidation and improved conformability Dhuley (2019). This enhancement relaxes the required surface finishing quality and contact pressure for copper interfaces.
@@
Another important component of IR shielding is an IR absorbing coating. In reference Bründermann and Hubers (2012), some absorbing coatings data, that are used in the IR radiation range, are provided (Eccosorb CR110, Herberts 1356H, TK RAM, and SiC grains in Stycast). Assuming that the IR shielding material does not transmit radiation, we can consider the sum of absorption and reflection coefficients to be 1. This allows us to estimate the absorption properties based on the reflection coefficient. According to theory Howell, Menguc, and Siegel (2010), absorbent coatings should exhibit maximum absorption (near 1) in the operating range. Among the absorbent coatings shown in reference Bründermann and Hubers (2012), SiC grains in Stycast have the highest absorption capacity. The literature demonstrates the prevalence of Stycast epoxy-based coatings for IR shielding. These coatings are employed both in pure form Grünhaupt et al. (2018); Maleeva et al. (2018) and when modified with additives to enhance surface roughness Barends et al. (2011) (Figure 3). For instance, mixtures of Stycast epoxy with carbon lampblack and 350 ${\mu}m$ SiC grit demonstrate absorption coefficients of 0.9 in the $THz$ frequency range. Barends et al. (2011). To create an optimal absorbing surface texture, various combinations Burnett et al. (2019); Byoung-moo (2021) of SiC particles of different sizes Dickel et al. (2018); Jin et al. (2015); Hazard (2019) and carbon powder Chou (2018); Wang (2022) may be utilized. Among these, the 1 $mm$ SiC particles dispersed in Stycast epoxy are particularly advantageous since their superior absorption performance. Such a specific particle size meets the surface roughness requirements for optimal IR absorption. Alternative absorber materials, such as Eccosorb resin epoxy Zhang (2018); Mundada (2021), may also be considered for these applications.
@@
In addition to IR protection, thin metallic layers are used to reflect spontaneous parasitic IR radiation. For this application, aluminized polyethylene terephthalate (commercially known as Mylar®) is typically employed Place et al. (2021); Gyenis et al. (2021); Narla (2017). The placement of this layer may vary: it may be placed around the holder Place et al. (2021), around the first shield Hazard (2019), between the shields Premkumar et al. (2021); Gyenis et al. (2021); Hazard (2019) or surround the entire shielding system Place et al. (2021); Premkumar et al. (2021); Hazard (2019). At the same time, the outer surface of the shield may act as a reflective layer due to its low absorption coefficient. This allows for the creation of a more compact shielding system.
@@
A more effective shielding system can be defined using the theory of a heat transfer by radiation and modeling the radiation propagation in the "sample-shielding-dilution refrigerator" system. A calculation scheme is represented in Figure 4a. In this system, "Body 1" is the sample, "Body 2" is the dilution refrigerator or radiation source, and "Shields" is the shielding system, which effectiveness we’d like to estimate. The power of a heat transfer by radiation is calculated by the following relation:
@@
In case the shield has different absorption coefficients of its inner and outer surfaces, formula for $A_{(1,2)sh}$ will be the following: $A_{(1,2)sh}=\frac{1}{\frac{1}{A_{1,2}}+\sum\limits_{i=1}^{n}\frac{F_{1}}{F_{shi}}\left({\frac{1}{A_{shi(inner)}}+\frac{1}{A_{shi(outer)}}-1}\right)};$
@@
Figure 4: a) Schematic of the radiative heat transfer model between two bodies with multiple shielding elements. b) Two-dimensional axisymmetric representation of the shielding system simulation. Key components: SQC (superconducting quantum circuit), SH (sample holder), FS (first shield), SS (second shield), C (cryostat’s lower-stage can), and HS (heat source).
@@
Figure 5: Schematic of the analyzed shielding configurations: A – uncoated copper sample holder with copper can attached to the lowest stage of a fridge; B –sample holder with absorption-coated can; C –sample holder with an inner absorption coating and uncoated can; D – sample holder with outer absorption coating and uncoated can; E – sample holder with both inner and outer coating and uncoated can; F – uncoated sample holder with inner coated first shield; G – inner coated sample holder with inner coated first shield and uncoated can; H – inner coated sample holder with two inner coated shields and uncoated can.
@@
The results of the calculations is verified using finite element simulations. Computer calculations is performed using COMSOL Multiphysics Heat Transfer in Solids and Surface to Surface Radiation. A two-dimensional axisymmetric model (see Figure 4b) was constructed to simplify calculations.In Figure 4b we introduce the following designations: “SQC” – superconducting quantum circuit, “SH” – sample holder, “FS” – first shield, “SS” – second shield, “HS” – heat source in the form of an IR filter, and “C” – can on the lowest stage of the fridge.
In our model, we utilize identical surface absorption coefficients for all computational scenarios. The absorption coefficient for silicon ($A_{Si}$) is 0.00008 Wollack et al. (2020), while for aluminum ($A_{Al}$) it is 0.0012 Benbow and Lynch (1975). The copper coefficient ($A_{Cu}$) is set to 0.005 Biondi (1956), and the absorptive coating ($A_{Absorp.coat.}$) has a value of 0.9 Barends et al. (2011). The system operates with a fixed heat dissipation power of $1\cdot 10^{-14}W$ (this evaluation has been done on the base of heat loads calculations carried out in Ref.Krinner et al. (2019)). The temperature of the cryostat’s lower stage is maintained at 0.01 $K$ throughout the simulations. We examine two distinct thermal scenarios: one where the superconducting quantum circuit itself serves as the heat source (designated SQC in Figure 4b), and another where an external heat source (designated HS in Figure 4b) provides the thermal input.